# Variational Mode-Driven Graph Convolutional Network for Spatiotemporal Traffic Forecasting

### v1

@@ III-B Architecture Detail
The architecture of the framework, as illustrated in Fig. 3, consists primarily of two steps: pre-processing and learning. In the initial stage of this architecture, the time series signal of each node is decomposed into $K$ modes or components. The number of modes $K$ is a hyper-parameter that must be adjusted based on the nature of the dataset. After recursively decomposing the signal, the decomposed components are concatenated along with additional features $d^{{}^{\prime}}$ such as time of the day and day of the week. The new features of the graph is formed into $\tilde{\mathcal{X}}\in\mathbb{R}^{N\times(K+d^{{}^{\prime}})\times T}$.
@@
The spatio-temporal (ST) block comprises spatial and temporal attention, graph convolution, and time convolution. Spatial and temporal attention mechanisms are applied to capture spatio-temporal correlations. Graph neural networks (GNNs) analyze the topological attributes of the graph and Chebyshev graph convolution (ChebNet) captures the properties of the graph for the Laplacian matrix in the spectral domain. Time convolution is employed to capture the temporal relationships within the graph structure. A residual connection combines the original feature with the output of the ST block, and 2D convolution is applied to its outcome. Multiple ST blocks are stacked together to obtain the final predicted outcome from the model.
IV Experiments
We perform our analysis for the following three case scenarios
@@ IV Experiments
  • VMGCN-v1: Features include original signal, modes, and time information.
@@
We compute the mean absolute percentage error (MAPE), mean absolute error (MAE), and root mean square error (RMSE) metrics to evaluate the performance of the proposed network. Before analyzing the performance of our proposed model, we will first provide a brief description of the dataset and the baseline models.
IV-A Datasets and Implementation Details
@@
In this work, for evaluation purposes, LargeST [36] dataset is used. It consists of a total of $8600$ nodes that are divided into four categories based on location: Greater Los Angeles (GLA), Greater Bay Areas (GBA), San Diego (SD), and others. It contains the traffic counts with a time interval of 5 minutes from 2017-2021. It also comprises meta-features such as longitude, latitude, and number of lanes.
@@
TABLE I: Distribution of nodes in LargeST where B (billion): $10^{9}$. Edges: Total number of edges between the nodes in the region, Degree: the average degree in each node.
@@
IV-B Baseline Models
  • Historical last (HL) [40] is a naive method that makes predictions based on the last observation.
  • Long short-term memory (LSTM) [41], is a variant of RNN to capture large temporal dependencies.
  • Diffusion convolutional recurrent neural networks (DCRNN) [12] captures the spatial and temporal pattern using diffusion convolution and encoding-decoding architecture respectively.
  • Adaptive graph convolutional recurrent networks (AGCRN) [7] determines the node-specific spatial and temporal correlation in traffic series.
  • Graph WaveNet (GWNET) [6] is a GNN architecture for spatial-temporal modeling.
  • Spatial-temporal graph ode networks (STGODE) [10] extracts the dynamical pattern from traffic data using ordinary differential equation (ODE).
  • Dynamic spatial-temporal aware graph neural networks (DSTAGNN) [19] constructs the data-driven graph and represents the spatial relevance using multi-head attention mechanism and temporal via gated convolution.
  • Dynamic graph convolutional recurrent networks (DGCRN) [9] employs a mechanism to generate the dynamical topology of a graph for traffic prediction.
  • Decoupled dynamic spatial-temporal graph neural networks (D2STGNN) [8] incorporates the decoupling of the traffic data in the learning mechanism.
  • Attention-based spatial-temporal graph convolutional networks (ASTGCN) [11] introduces the attention mechanism and applies the graph convolution and temporal convolution in the modeling of ST data.
IV-C Analysis
@@
(b) $\sum\limits_{i=0}^{8}\lvert\hat{u_{i}}(\omega)\rvert$
@@
Figure 4: Traffic signal decomposition. a) The normalized traffic signal of a sensor b) the spectrum of the magnitude of the reconstructed signal from modes. c) IMF1 - IMF8 plotted from lowest to highest-frequency components from top to bottom, left to right order.
VMD recursively decomposes traffic data into modes that can be categorized into high-frequency, intermediate, and low-frequency components. These modes represent the real-time non-stationary and complex nature of the signals, which consist of numerous modes, as depicted in Fig. 4. Fig. 4(a) displays the counts from a 1-year traffic sensor in the SD region. The pattern of periodic, quasi-periodic, and noise components in the signal spectrum is illustrated in Fig. 4(b). The signal is segregated into 8 modes, shown in Fig. 4(c) ordered from low to high frequencies (IMF1 - IMF8). These modes can have physical interpretation in real-time traffic analysis. The lowest-frequency mode (IMF1) mimics the overall trend of the original signal at low frequency, while the highest-frequency mode (IMF8) describes the noise in the signal. The number of intermediate-frequency modes may vary depending on the total number of modes and the complexity of the signal. The noise-free signal can be reconstructed by summing the components, excluding the lowest-frequency and highest-frequency modes, as referenced in [31].
The tuning of hyper-parameters remains a crucial step in the training process of a deep learning model. In the decomposition of data, it is essential to select the appropriate values for the number of modes ($K$), bandwidth constraint ($\alpha$), tolerance for convergence ($\epsilon$), and total iteration ($N_{H}$) for VMD. The performance of our proposed method has been evaluated under different values of $K$ modes. Fig. 5 shows the MAE training loss evaluated on the SD region of the LargeST dataset, while keeping the other hyper-parameters constant. It shows that the loss gradually decreases as the number of modes increases. However, finding the suitable value of $K$ that yields the best prediction results is computationally expensive. Consequently, we have formulated a method that utilizes reconstruction loss based on the threshold ($\zeta$) criteria.
@@
Figure 5: Training loss with different modes on SD dataset for VMGCN-v1. It shows that the training loss decreases with the increase in the number of modes.
To determine the best-fit number of mode, we randomly select the nodes and compute the modes and for each, subsequently calculating the average reconstruction loss of the sample data for every $K$ value. If the reconstruction loss is less than the threshold value ($\zeta$), we can consider this $K$ value as a near-best term. In our experimentation, we selected $2\%$ of the nodes from each region and applied this method for $K$ values ranging from $2$ to $29$, as shown in Fig. 6. The best value of $K$ is identified for each region: SD $K=13$, GBA $K=14$, GLA $K=13$ when the reconstruction loss is less than the $\zeta=1\times 10^{-3}$. After identifying the $K$ value for each region, it is observed that any further decrease in loss beyond these values is negligible.
@@
TABLE II: Comparison of performance evaluation metrics MAE, MAPE, and RMSE between different advanced methods and our models. The number of parameters (param) is described in K (kilo), $10^{3}$ and M (million), $10^{6}$, and the best performance metrics are highlighted in red bold numbers. For VMGCN-vx-$K$, x describes the variant of the network and $K$ corresponds to the best value chosen for each region.
@@
The comparison of our experimentation with baseline models is presented in TABLE II. We evaluated our case scenarios using $\alpha=2000$, $\epsilon=1\times 10^{-7}$, and $N_{H}=500$. For VMGCN-v1, incorporating the original feature with the modes generally leads to better predictions. In long-horizon prediction, VMGCN-v3 exhibits relatively lower error values compared to VMGCN-v2. The decomposition by VMD is sensitive to the length of the signal. Currently, utilizing the 1-year data makes the computed modes more effective. Typical traffic trends such as jams, congestion, and daily, or weekly patterns are well-defined in terms of frequencies. Due to the accurate representation of features, the neural networks predict future states. Increasing the number of modes enhances the performance of the model in short-term prediction, but it tends to overfit in long-term predictions. It appears that VMD effectively decomposes the signal at higher modes and removes noise, leading to more precise immediate prediction.
Figure 6: The sub-optimal number of modes is determined by randomly selecting signal and performing decomposition for mode range $[2,29]$. If the mean reconstruction loss of the signals is less than a threshold value on a specific mode, consider that mode as a sub-optimal value.
(a) MAE
(b) RMSE
(c) MAPE
Figure 7: The importance of each mode in horizon prediction is studied by replacing the respective mode with zero value and performing inference to observe the change in the performance metrics. The plot shows the differences in metrics between the original and distorted modes on the logarithmic scale. The greater the difference, the more significant the mode for that horizon. This analysis has been performed on the SD region for VMGCN-v2.
Figure 8: This plot shows the comparison of ground truth with the predicted traffic flow for horizon 1 on three distinct locations in the SD region.
To evaluate the effectiveness of each mode on horizon prediction, the modes are analyzed in terms of central frequencies and energy. This relationship between prediction accuracy and the modes, separated based on frequencies, is investigated in Fig. 7. In VMGCN-v2 for the SD region, inference is performed by setting each mode to zero in the test set. The Fig. 7(a)7(b), and 7(c) show the logarithm of non-zero $\Delta$ MAE, $\Delta$ RMSE, and $\Delta$ MAPE repeativety for all horizons, respectively. Here, $\Delta$ represents the difference between the metric with all features and the metric with zero value in $k^{\rm{th}}$ mode for a horizon. The modes from $1$ to $13$ are arranged in terms of increasing frequency components. A higher metric value signifies the importance of that mode in horizon prediction. It has been observed that lower-frequency components contribute more to prediction accuracy relative to high-frequency components. As frequency increases, the noise in the mode also increases, and the pattern useful for prediction decreases. Removing the high-frequency components during the training of the neural network could potentially enhance the capabilities of the model performance.
We also compare the predicted traffic flow and the actual flow for Horizon $1$ plotted in Fig. 8 for three sensors at different locations. The model under-predicts some values at the peak times, which correspond to rush hours. Sensors 2 and 3 exhibit some periodic trends with harmonic noise.
TABLE III: Performance evaluation of metrics MAE, MAPE, and RMSE
@@
IV-D Ablative Study

### v3

@@ 3.2. Model Architecture
The model architecture, as illustrated in Fig. 3, consists of two steps: pre-processing and learning. In the first stage of this architecture, the time series signal of each node is decomposed into $K$ modes or components. The number of modes $K$ is a hyper-parameter that must be adjusted based on the nature of the dataset. After recursively decomposing the signal, the decomposed components are concatenated along with $d^{{}^{\prime}}$ number of additional features such as time of the day and day of the week. The new features of the graph are formed into $\mathcal{X}\in\mathbb{R}^{N\times(K+d^{{}^{\prime}})\times T_{w}}$.
@@
The spatiotemporal (ST) block comprises spatial and temporal attention, graph convolution, and time convolution. Spatial and temporal attention mechanisms are applied to capture spatiotemporal correlations. Graph neural networks (GNNs) analyze the topological attributes of the graph, and Chebyshev graph convolution (ChebNet) captures the properties of the graph for the Laplacian matrix in the spectral domain. It is essential to capture global dependencies to incorporate the structural information of a road network, as illustrated in Fig. 4. By increasing the order $M$ of the Chebyshev polynomial, the model aggregates information from nodes up to $M$ hops away. Time convolution is employed to capture the temporal relationships within the graph structure. A residual connection combines the original feature with the output of the ST block, and a 2D convolution is applied to its outcome. Multiple ST blocks are stacked together to obtain the final predicted outcome from the model.
Figure 4. Left picture depicts the mechanism to count the vehicles at each intersection. The right picture explains the local and global dependencies, where global dependencies are learned through a graph network in each hop.
4. Experiments
We perform our analysis for the following three cases:
@@ 4. Experiments
  • VMGCN-v1: Features include original signal, modes, and time information (time of day, day of week).
@@
We compute the mean absolute percentage error (MAPE), mean absolute error (MAE), and root mean square error (RMSE) metrics to evaluate the performance of the proposed network. Before analyzing the performance of our proposed model, we first provide a brief description of the dataset and the baseline models.
4.1. Datasets and Implementation Details
@@
We use LargeST (Liu et al., 2024) dataset for comparing the performance of VMGCN with the existing models. It consists of a total of $8600$ nodes that are divided into four categories based on location: Greater Los Angeles (GLA), Greater Bay Areas (GBA), San Diego (SD), and others. It contains the traffic counts with a time interval of 5 minutes from 2017-2021. It also comprises meta-features such as longitude, latitude, and number of lanes. We summarize the number of edges, nodes, the average degree of each node, and the total data points (in billions, B) in the LargeST dataset in Table I.
@@
Table 1. LargeST dataset summary.
@@
4.2. Baseline Models
  • Historical last (HL) (Liang et al., 2021) is a naive method that makes predictions based on the last observation.
  • Long short-term memory (LSTM) (Hochreiter and Schmidhuber, 1997), is a variant of RNN to capture large temporal dependencies.
  • Diffusion convolutional recurrent neural networks (DCRNN) (Li et al., 2018) captures the spatial and temporal pattern using diffusion convolution and encoding-decoding architecture, respectively.
  • Adaptive graph convolutional recurrent networks (AGCRN) (Bai et al., 2020) determines the node-specific spatial and temporal correlation in traffic series.
  • Graph WaveNet (GWNET) (Wu et al., 2019) is a GNN architecture for spatial-temporal modeling.
  • Spatial-temporal graph ode networks (STGODE) (Fang et al., 2021) extracts the dynamical pattern from traffic data using ordinary differential equation (ODE).
  • Dynamic spatial-temporal aware graph neural networks (DSTAGNN) (Lan et al., 2022) constructs the data-driven graph and represents the spatial relevance using multi-head attention mechanism and temporal via gated convolution.
  • Dynamic graph convolutional recurrent networks (DGCRN) (Li et al., 2023) employs a mechanism to generate the dynamical topology of a graph for traffic prediction.
  • Decoupled dynamic spatial-temporal graph neural networks (D2STGNN) (Shao et al., 2022) incorporates the decoupling of the traffic data in the learning mechanism.
  • Attention-based spatial-temporal graph convolutional networks (ASTGCN) (Guo et al., 2019) introduces the attention mechanism and applies the graph convolution and temporal convolution in the modeling of ST data.
  • Spatio-Temporal Wavelets (STWave) (Fang et al., 2023) employed the discrete wavelet transform (DWT) to convert the traffic time series into low-frequency and multi-high-frequency components, then use their inverse DWT (IDWT) representation. Dual-channel ST encoder uses temporal attention and causal convolution to learn time dependencies, while spectral graph attention networks have been incorporated to capture spatial dynamics. And at the end, two decoders forecast the future traffic trends.
  • BigST (Han et al., 2024) pre-processes the long-term historical features by using a generative pre-training method. This framework also quantifies the spatial dependencies using attention scores to construct the graph structures.
  • Random projection mixer (RPMixer) (Yeh et al., 2024) converts the time series data into the frequency domain using Fast Fourier Transform (FFT) and passes each real and imaginary component to a linear layer. Random projection layers have been used to enhance the spatial modeling.
  • Patch spatio-temporal graph (PatchSTG) (Fang et al., 2025) implements the irregular spatial patching by utilizing the leaf K-dimensional tree (KDTree) to capture the spatial knowledge dynamically.
  • Regularized adaptive graph learning (RAGL) (Wu et al., 2025) achieves embedded regularization and noise suppression by introducing the stochastic shared embedding and an adaptive graph convolution (AGC) encoder in their framework to forecast the future states.
4.3. Analysis
@@
(b) $\sum\limits_{i=1}^{8}\lvert\hat{u_{i}}(\omega)\rvert$
@@
(d) Reconstructed signal (weekday)
(e) Reconstructed signal (weekend)
[…]
Variational Mode Decomposition (VMD) recursively decomposes traffic data into modes that can be categorized as high-frequency, intermediate-frequency, and low-frequency components. These modes capture the real-time, non-stationary, and complex nature of traffic signals, which typically consist of multiple overlapping patterns, as shown in Fig. 5. Fig. 5(a) displays the counts from a 1-year (on a scale 0 to 1) traffic sensor in the SD region. The signal spectrum, illustrated in Fig. 5(b), reveals the presence of periodic, quasi-periodic, and noise components. The signal is decomposed into eight Intrinsic Mode Functions (IMFs), labeled IMF1 to IMF8 and ordered from low to high frequency, as shown in Fig. 5(c). These modes allow for physical interpretation in the context of real-time traffic analysis. IMF1 represents the overall trend of the original signal, capturing its low-frequency behavior, while IMF8 captures the high-frequency elements or noise. The intermediate modes (IMF2 to IMF7) capture various degrees of temporal patterns depending on the complexity of the original signal. A denoised version of the signal can be reconstructed by summing selected IMFs, typically excluding the lowest and highest frequency components, as outlined in (Dragomiretskiy and Zosso, 2013). Fig. 5(d) and Fig. 5(e) provide a detailed view of the decomposed modes and the reconstructed signals over a 24-hour period on a weekday (Monday) and a weekend (Saturday), respectively. These plots show that IMF1 acts as the baseline trend, while IMF2 and IMF3 capture the envelope and overall structure of the signal. Higher-frequency IMFs account for short-term fluctuations. Notably, during early morning hours (1:00 AM to 3:00 AM), traffic volume is significantly lower on weekends compared to weekdays. This pattern has been successfully captured by IMF2 and IMF3. Each mode effectively represents aspects of traffic flow and human activity patterns, which can ultimately enhance the ability of neural networks to learn trends and make accurate traffic flow predictions.
Figure 6. Index of orthogonality on a traffic signal of the SD region. (a) shows the variation in IO with different $\alpha$ and $K$. Visualization of (b) $\sum_{t=0}^{T}u_{i}(t)\,u_{j}(t)$, and (c) $\frac{\sum_{t=0}^{T}u_{i}(t)\,u_{j}(t)}{\sum_{t=0}^{T}f^{2}(t)}$ among decomposed components for $K=8$ and $\alpha=2000$.
The tuning of hyperparameters remains a crucial step in solving the optimization process. In the decomposition of data, it is essential to select the appropriate values for the number of modes ($K$), bandwidth constraint ($\alpha$), tolerance for convergence ($\epsilon$), and maximum allowable iterations ($N_{m}$) for VMD. In this work, we consider three measures to explore the significance of VMD hyperparameters. The index of orthogonality (IO) provides a quantitative measure of the temporal independence between decomposed modes (Huang et al., 1998). A lower IO indicates that the modes are non-orthogonal, better mode separation, minimal redundancy, and improved interpretability, making it an essential quality metric in signal decomposition. Ideally, this metric must be zero for off-diagonal values, but in a practical sense, it would be a non-zero, and it can be measured by
Figure 7. Spectral overlap index computed on the decomposed modes of a traffic signal of the SD region. (a) $K=2,\,\alpha=20$, (b) $K=13,\,\alpha=20$, (c) $K=2,\,\alpha=1000$, (d) $K=13,\,\alpha=1000$, (e) $K=2,\,\alpha=2000$, and (f) $K=13,\,\alpha=2000$.
(14) $\text{IO}=\left(\frac{\sum\limits_{i=1}^{K+1}\sum\limits_{j=1}^{K+1}\sum\limits_{t=0}^{T}u_{i}(t)\,u_{j}(t)}{\sum\limits_{t=0}^{T}f^{2}(t)}\right),$
where the additional component indicates the redemption or residual function $\phi(t)$. Fig. 6(a) shows the trend of IO with different bandwidth constraint values ($\alpha$) and number of modes ($K$). Overall, the IO decreases with increasing value of the number of modes with higher $\alpha$ values. However, with a lower value of $\alpha=20$, the IO values keep increasing. Increasing the $\alpha$ decreases the bandwidth and vice versa. For a higher number of modes and a lower value of the bandwidth constraint, the overlaps occur in modes. Higher value of bandwidth constraint and number of modes avoids overlapping because of narrow bandwidth. The average value of IO for $20$ nodes of the SD region is $0.0661$ for $K=13$ and $\alpha=2000$. Fig. 6(b) represents the unnormalized heatmap $\sum\limits_{t=0}^{T}u_{i}(t)\,u_{j}(t)$ and Fig. 6(c) illustrates the normalized IOi,j for $K=8$ and $\alpha=2000$ between the modes. It shows that off-diagonal entities are zero and three components (IMF1-IMF3) share the energy, having a cumulative value of $0.95$, and the remaining components contribute to $0.05$ energy of the total signal. The spectral overlap index (SOI) measures the shared frequency content between decomposed modes. Lower SOI values reflect better spectral separation, indicating that each mode captures a distinct frequency band, essential for interpretable results and can be formulated as
(15) $\text{SOI}_{i,j}=\frac{\sum_{\omega}\hat{u}_{i}(\omega)\,\hat{u}_{j}(\omega)}{\sqrt{\sum_{\omega}\hat{u}_{i}^{2}(\omega)\sum_{\omega}\hat{u}_{j}^{2}(\omega)}},$
here $i,j\in[1,K+1]$ includes the residual component in frequency domain $\Phi(\omega)$. By increasing $\alpha$, the modes become more spectrally distinct, as indicated in Fig. 7. This is further evidenced by the decrease in the SOI observed in Fig. 7(e) and Fig. 7(f), compared to Fig. 7(a) and Fig. 7(b), respectively. The difference between Fig. 7(e) and Fig. 7(c) or Fig. 7(f) and Fig. 7(d) is negligible due to higher value of $\alpha$. As we mentioned previously, increasing the number of modes decreases the deformation value; the same trend can be observed in measuring SOI. The frequency overlapping between the decomposed modes and the residual component decreases as $K$ and $\alpha$ increase. The residual is also important to determine suitable hyperparameters, as it is noticed that increasing $K$ and decreasing $\alpha$ results in low reconstruction loss. The small reconstruction loss preserves the information loss from the original signal during decomposition. To determine the optimal value of number of modes, we randomly select nodes, compute the modes for each, and subsequently calculate the average reconstruction loss of the sample data for every $K$ value. If the reconstruction loss is less than the threshold value ($\zeta$), we can consider this $K$ value as a near-best term. In our experimentation, we selected $2\%$ of the nodes from each region and applied this method to values of $K$ ranging from $2$ to $29$, as shown in Fig. 8. The optimal value of $K$ is identified for each region: SD ($K=13$), GBA ($K=14$), and GLA ($K=13$), for which the reconstruction loss remains below the threshold $\zeta=1\times 10^{-3}$. After identifying the $K$ value for each region, it is observed that any further decrease in loss beyond these values is negligible. It must be noted that our reconstruction-loss-based mode selection (Fig.  8) eliminates heuristic tuning that is a common pitfall in EMD-based methods. This data-driven selection of $K$ ensures adaptability across regions. The residual component $\phi(t)$ and bandwidth constraint $\alpha$ explicitly filter sensor noise. This contrasts with wavelet-based methods (e.g., STWave) where noise may alias into multiple frequency bands.
Figure 8. The optimal value of $K$ is determined by randomly choosing 2% of the node signals in a region of interest and decomposing these for $K\in[2,29]$. If the mean reconstruction loss is less than a threshold value for a specific value of the mode, we consider that mode as an optimal value of $K$.
The performance of our proposed method has been evaluated under different values of $K$ modes. Fig. 9 shows the MAE training loss evaluated on the SD region of the LargeST dataset, while keeping the other hyper-parameters constant. It shows that the loss gradually decreases as the number of modes increases. However, finding a suitable value of $K$ that yields the best prediction results is computationally expensive. Consequently, we have formulated a method that utilizes the reconstruction loss to determine the optimal value of $K$. The comparison of our experimentation with baseline models is presented in TABLE 2. Our model outperforms the signal processing-based decomposition methods that utilize the Wavelet transform and Fourier transform, such as STWave and RPMixer, and representation learning based decomposition methods such as D2STGNN. We evaluated our case scenarios using $\alpha=2000$, $\epsilon=1\times 10^{-7}$, and $N_{m}=500$. For VMGCN-v1, incorporating the original feature with the modes generally leads to better predictions. In long-horizon prediction, VMGCN-v3 exhibits relatively lower error values compared to VMGCN-v2. The decomposition by VMD is sensitive to the length of the signal. Currently, utilizing the 1-year data makes the computed modes more effective. Typical traffic trends such as jams, congestion, and daily, or weekly patterns are well-defined in terms of frequencies. Due to the accurate representation of features, the neural networks predict future states. Increasing the number of modes enhances the performance of the model in short-term prediction, but it tends to overfit in long-term predictions. It appears that VMD effectively decomposes the signal at higher modes and removes noise, leading to more precise immediate prediction.
@@
Figure 9. Training loss with different modes on SD dataset for VMGCN-v1. It shows that the training loss decreases with the increase in the number of modes.
@@
Table 2. Comparison of performance evaluation metrics MAE, MAPE, and RMSE between different advanced methods and our models. The number of parameters (param) is described in K (kilo), $10^{3}$ and M (million), $10^{6}$, and the best performance metrics are highlighted in red bold numbers. For VMGCN-vx-$K$, x describes the variant of the network and $K$ corresponds to the best value chosen for each region. × parameters count is not publicly available.
@@
    results executed by (Fang et al., 2025).
To evaluate the effectiveness of each mode on horizon prediction, the modes are analyzed in terms of central frequencies and energy. This relationship between prediction accuracy and the modes, separated based on frequencies, is investigated in Fig. 10. In VMGCN-v2 for the SD region, inference is performed by setting each mode to zero in the test set. Fig. 10(a), Fig. 10(b), and Fig. 10(c) show the logarithm of non-zero $\Delta$ MAE, $\Delta$ RMSE, and $\Delta$ MAPE for all horizons, respectively. Here, $\Delta$ represents the difference between the metric with all features and the metric with zero value in $k^{\rm{th}}$ mode for a horizon. The modes from $1$ to $13$ are arranged in terms of increasing frequency components. A higher metric value signifies the importance of that mode in horizon prediction. It has been observed that lower-frequency components (IMF1-IMF4) contribute more to prediction accuracy relative to high-frequency components. As frequency increases, the noise in the mode also increases, and the pattern useful for prediction decreases. High-frequency components (IMF7-IMF13) are significant in short-term predictions; their contributions to long-term predictions decrease. Removing the high-frequency components during the training of the neural network could potentially enhance the performance capabilities of the model. We also compare the predicted traffic flow and the actual flow for Horizon $1$ plotted in Fig. 11 for three sensors at different locations. It has been observed that the model under-predicts some values at peak times that correspond to rush hours, and sensors 2 and 3 exhibit some periodic trends with harmonic noise.
Table 3. Analysis of performance evaluation metrics MAE, MAPE, and RMSE for different VMD parameters (SD region).
@@
(a) MAE
(b) RMSE
(c) MAPE
Figure 10. The importance of each mode in horizon prediction is studied by replacing the respective mode with a zero value and performing inference to observe the change in the performance metrics. The plot shows the difference in metrics between the original and distorted modes on the logarithmic scale. The greater the difference, the more significant the mode for that horizon. This analysis has been performed on the SD region for VMGCN-v2.
4.4. Ablative Study